the perimeter of a triangle is 132cm and the lengths of its sides are in the ratio 8:11:14 . find the length each side ?
Answer by mananth(16946)

perimeter =132
proportion = 8:11:14
total parts 33
8/33*132=32 one side
..
11/33*132=44 second side
..
14/33*132=56 third side.
..